# Reply with Sticker: New Dataset and Model for Sticker Retrieval

### v4

@@ Abstract
Using stickers in online chatting is very prevalent on social media platforms, where the stickers used in the conversation can express someone’s intention/emotion/attitude in a vivid, tactful, and intuitive way. Existing sticker retrieval research typically retrieves stickers based on context and the current utterance delivered by the user. That is, the stickers serve as a supplement to the current utterance. However, in the real-world scenario, using stickers to express what we want to say rather than as a supplement to our words only is also important. Therefore, in this paper, we create a new dataset for sticker retrieval in conversation, called StickerInt, where stickers are used to reply to previous conversations or supplement our words. Based on the created dataset, we present a simple yet effective framework for sticker retrieval in conversation based on the learning of intention and the cross-modal relationships between conversation context and stickers, coined as Int-RA. Specifically, we first devise a knowledge-enhanced intention predictor to introduce the intention information into the conversation representations. Subsequently, a relation-aware sticker selector is devised to retrieve the response sticker via cross-modal relationships. Extensive experiments on two datasets show that the proposed model achieves state-of-the-art performance and generalization capability in sticker retrieval [^1] The dataset and source code of this work are released at https://github.com/HITSZ-HLT/Int-RA..
@@ I Introduction
Recent research endeavors of sticker retrieval have been dedicated to using stickers to supplement the current response, in order to strengthen the expression of emotion, attitude, or opinion [7, 8]. However, in the real-world scenario, we may also use stickers to reply to the previous conversation directly, rather than merely supplementing our words with stickers. We refer to the former as the Supplementary Response (SR) scenario, where stickers are used alongside the speaker’s own utterance, and to the latter as the Direct Response (DR) scenario, where stickers function as independent replies without accompanying text. One expected scenario of the sticker retrieval task is that suitable stickers can be retrieved for replies whether or not one has made a textual response. To illustrate our idea, we give examples shown in Figure 1. Figure 1 (a) shows the SR scenario, and Figure 1 (b) depicts the DR scenario. Therefore, in our work, we create a new sticker retrieval dataset to cover these two scenarios, called StickerInt, which is a comprehensive consideration for using stickers in social media conversation. Our StickerInt dataset contains 1,578 Chinese conversations with 11,066 utterances.
@@ II-B Image Retrieval
Image retrieval aims to retrieve images that match a given query from a large collection of images. Early method [11] mainly relied on global feature-based image-text matching. Further, [12] proposed a Visual Semantic Reasoning Network (VSRN), which generates global features with regional semantic associations through a Graph Convolutional Network (GCN). To address the false negative problem in existing image-text matching benchmarks, [13] proposed an Image-Text Matching (ITM) framework integrating Language Guidance. In terms of improving retrieval efficiency and accuracy, [14] proposed a self-supervised fine-grained alignment module called SelfAlign. [15] proposed a feature-first approach to advance image-text retrieval by improving visual features. This method generates more comprehensive and robust visual features through a two-step interaction model and a multi-attention pooling module. Recently, with the development of large-scale pretraining models like CLIP [16] and ALIGN [17], significant performance improvements have been achieved in ITM tasks by pretraining on massive vision-language data and then fine-tuning for specific tasks. Different from the image retrieval methods, we focus on retrieving the image based on the conversation not just the text.
@@ II-C Multi-modal Conversation
Several multi-modal studies have sought to improve the efficacy of conversational agents by enriching textual expressions of generated dialog responses through associative vision scenes [18, 19]. In contrast, sticker retrieval in the real-world requires understanding the semantic expression and intention of the context and the sticker image. Several works specifically paid attention to sticker features [20, 4]. [20] proposed a real-time system for sticker recommendation, which decomposes the sticker recommendation task into two steps, including message prediction that the user is likely to send and an appropriate sticker substitution. Nowadays, using stickers for replies has become commonplace in social media interactions. There has been a growing number of works on sticker retrieval, which assists users in selecting the appropriate sticker for a response. [8] proposed a sticker response selector model that employs a convolution-based sticker image encoder and a self-attention-based multi-turn dialog encoder to obtain the representation of stickers and utterances. Then, a deep interaction network is designed for obtaining deep matching, and a fusion network is employed to output the final matching score. [7] presented a Meme incorporated Open-domain Dialogue (MOD) task and utilized pooling all sub-tasks like text generation and internet meme prediction into a unified sequence generation procedure to solve it. To tackle three challenges of inherent multimodality, significant inter-series variations, and complex multi-modal sentiment fusion, [4] integrated both the sticker text and series tag to holistically model sticker sentiment. They employed a flexible masked attention mechanism to selectively extract the most pertinent information crucial for the current sentiment assessment. These methods, which directly match conversation context and stickers, overlook the expressive role of stickers in conversation, where emotions and intentions are conveyed in a visually engaging manner.
@@ III StickerInt Dataset
This section introduces our new dataset StickerInt for sticker retrieval in detail. Specifically, Section III-A presents the construction of our StickerInt dataset, Section III-B introduces the annotation process of StickerInt, and Section III-C and Section III-D analyze the created StickerInt dataset.
@@ III-C Dataset Analysis
The detailed statistics of our dataset are shown in Table I. In total, there are 1,578 conversations that contain 11,066 utterances, 59,623 tokens, and 1,025 stickers. Of these stickers, 702 are associated with textual content. The proposed dataset is collected from conversations with Chinese users on WeChat, and the dialogues are predominantly in Chinese. Notably, 841 utterances include English words such as “OK”, “QQ”, and “NLP”, which are naturally embedded in Chinese expressions, reflecting the way people commonly communicate in daily life. Each conversation includes 8.01 utterances on average. The average number of users who participate in a conversation is 2.80. Among all conversations in our dataset, 1,017 belong to the Supplementary Response scenario, and 561 fall under the Direct Response scenario (sticker as a direct reply to previous conversation turns). The average number of tokens in an utterance is 5.21.
Furthermore, we also visualize the distribution of intention tags in Figure 2. We can see that the proportion of stickers used in different labels varies, which also demonstrates the diversity of stickers’ intention expressions in our dataset.
@@ IV Methodology
In this section, we introduce our proposed Int-RA framework for sticker retrieval in detail. Assume that there is a conversation context $T=\{t_{1},...,t_{N_{t}},s_{j}\}$ and a sticker set $V=\{v_{1},...v_{N_{v}}\}$, where $t_{i}=(s_{i},u_{i})$, $s_{i}$ and $u_{i}$ represent the $i$-th user and utterance. $N_{t}$ and $N_{v}$ represent the number of utterances and stickers, respectively. The sticker retrieval task aims to select a suitable sticker from the sticker set $V$ based on the conversation context $T$ for the user $s_{j}$. In the $i$-th utterance $u_{i}=\{w_{1}^{i},...,w^{i}_{N_{w}^{i}}\}$, $w_{j}^{i}$ represents the $j$-th word in $u_{i}$, and $N_{w}^{i}$ represents the total number of words in $u_{i}$ utterance. Additionally, our dataset presents an annotated intention tag $y_{int}$ as supplemental information in each conversation $T$. In the dialog context $T$, $v_{i}$ represents a sticker image with a binary label $y_{i}$, indicating whether $s_{i}$ is an appropriate response for $T$. Each sticker with the intention label $m_{i}$ indicates the intention of the speaker to use the sticker to respond. Our goal is to learn a ranking model that can produce the correct ranking for each dialog.
@@ V-C Main Results
For text-to-image retrieval methods, they prioritize capturing semantic relationships between text and image content. However, since they are not explicitly designed for sticker retrieval scenarios, they exhibited inferior performance compared to our framework. In addition, an interesting observation is the overall superior performance of text-to-image retrieval methods compared to sticker retrieval baseline methods. This disparity can be attributed in part to the model design of SRS and MOD, where both models are devised to retrieve suitable stickers from a limited set of similar sticker candidates. Consequently, more attention is devoted to distinguishing local information among similar sticker expressions. In contrast, our dataset encompasses a more diverse range of stickers in real-world conversations, requiring the model to analyze and recognize more complex visual features. This demonstrates the advantage of the new dataset and the proposed framework in this work.
For large language models, except for Ultron and ChatIR, we design a prompt that integrates the current session to generate the sticker description for each session, leveraging the in-context learning ability of LLMs [47, 48]. We then retrieve the appropriate sticker based on the generated sticker description and the combination of intent and sticker attributes using OpenAI’s LLM-based embedding model (text-embedding-ada-003) [49]. We can observe that LLM models perform better than sticker retrieval and text-to-image retrieval methods. This superior performance is attributed to the extensive parameterization and complex network architecture of LLMs, which greatly enhance their ability to understand and generate complex language and image descriptions. For P@1, Ultron performs the best, while for P@3 and P@5, ChatGLM3 achieves the highest performance, reaching 25.32% and 31.17%. Compared to our Int-RA method, all baseline models perform significantly worse. This further emphasizes the effectiveness of our approach in intent derivation, highlighting the critical role of intent as a bridging component in the process.
@@ V-E Ablation Study
We also conduct an ablation study on the use of knowledge and attributes. The evaluation results are shown in Table IV. The performances of all ablation models are worse than those of the complete model under all metrics, which demonstrates the necessity of each component in our approach. Note that the removal of attributes ("w/o attribute") results in considerable performance degradation, indicating that utilizing attributes can lead to better learning of sticker representation in different sticker properties. Notably, by observing the performance of w/o attribute on P@5, we find that the impact of attributes is not as significant when a larger number of stickers are recalled. In addition, the removal of the intention tag ( "w/o intention" ) sharply degrades performance, which verifies the importance of knowledge in understanding conversation context. This demonstrates the crucial role of the intention tag in improving the accuracy and relevance of sticker retrieval tasks, highlighting the importance of considering users’ intentions in retrieving appropriate stickers.
@@ V-G Effect of Number of Utterances
To examine and analyze the impact of the number of utterances on the performance of our proposed Int-RA framework, we conduct experiments by varying the number from 2 to 10 and demonstrate the results in Figure 6. We observe a similar trend across all evaluation metrics: mAP, P@1, P@3, and P@5. The results initially increase until the number of utterances reaches 6, after which they decrease as the number of utterances continues to increase. One potential reason may explain this phenomenon. While more context (up to six utterances) enriches the model’s understanding and improves performance, excessive context introduces noise from irrelevant earlier turns, impairing accuracy. Our findings show that a six-utterance context strikes the optimal balance between information richness and noise, leading to the best performance for sticker selection.
@@ V-K Visualization
To analyze how our Int-RA learns the important information about stickers, we visualize the relation score $\mathcal{M}$ (Equation 12) of three stickers in Figure 8. For example (a), where the character appears neutral. This indicates that the representation of this sticker heavily relies on this facial expression. Our Int-RA can effectively catch the important information in the sticker by the relation score placement on the character’s face. Moreover, the relation score also attends to the character’s gestures. For instance, in example (b), where the character is depicted as holding chopsticks with one hand and supporting the face with the other, we observe attention focused on his hand, suggesting that our Int-RA learns key points of body language. Furthermore, considering that the relation score comprehensively considers four properties of stickers, as illustrated in example (c), we observe that our Int-RA pays attention to both facial expressions and gestures simultaneously, thereby learning accurate visual information about the sticker.
@@ Acknowledgments
This work was partially supported by the National Natural Science Foundation of China (62176076), the Natural Science Foundation of Guangdong (2023A1515012922), the Shenzhen Foundational Research Funding (JCYJ20220818102415032), the Major Key Project of PCL2023A09, Hong Kong RGC GRF No. 14206324, CUHK direct grant No. 4055209, and CUHK Knowledge Transfer Project Fund No. KPF23GWP20.

### v5

@@ Abstract
Using stickers in online chatting is very prevalent on social media platforms, where the stickers used in the conversation can express someone’s intention/emotion/attitude in a vivid, tactful, and intuitive way. Existing sticker retrieval research typically retrieves stickers based on context and the current utterance delivered by the user. That is, the stickers serve as a supplement to the current utterance. However, in the real-world scenario, using stickers to express what we want to say rather than as a supplement to our words only is also important. Therefore, in this paper, we create a new dataset for sticker retrieval in conversation, called StickerInt, where stickers are used to reply to previous conversations or supplement our words [^1] We believe that the release of this dataset will provide a more complete paradigm than existing work for the research of sticker retrieval in the open-domain online conversation.. Based on the created dataset, we present a simple yet effective framework for sticker retrieval in conversation based on the learning of intention and the cross-modal relationships between conversation context and stickers, coined as Int-RA. Specifically, we first devise a knowledge-enhanced intention predictor to introduce the intention information into the conversation representations. Subsequently, a relation-aware sticker selector is devised to retrieve the response sticker via cross-modal relationships. Extensive experiments on two datasets show that the proposed model achieves state-of-the-art performance and generalization capability in sticker retrieval [^2] The dataset and source code of this work are released at https://github.com/HITSZ-HLT/Int-RA..
@@ I Introduction
Recent research endeavors of sticker retrieval have been dedicated to using stickers to supplement the current response, in order to strengthen the expression of emotion, attitude, or opinion [7, 8]. However, in the real-world scenario, we may also use stickers to reply to the previous conversation directly, rather than merely supplementing our words with stickers. We refer to the former as the Supplementary Response (SR) scenario, where stickers are used alongside the speaker’s own utterance, and to the latter as the Direct Response (DR) scenario, where stickers function as independent replies without accompanying text. One expected scenario of the sticker retrieval task is that suitable stickers can be retrieved for replies whether or not one has made a textual response. To illustrate our idea, we give examples shown in Figure 1. Figure 1 (a) shows the SR scenario, and Figure 1 (b) depicts the DR scenario. Therefore, in our work, we create a new sticker retrieval dataset to cover these two scenarios, called StickerInt, which is a comprehensive consideration for using stickers in social media conversation. Our StickerInt dataset contains 1,578 Chinese conversations with 12,644 utterances.
@@ II-B Image Retrieval
Image retrieval aims to retrieve images that match a given query from a large collection of images. Early method [11] mainly relied on global feature-based image-text matching. Further, [12] proposed a Visual Semantic Reasoning Network (VSRN), which generates global features with regional semantic associations through a Graph Convolutional Network (GCN). To address the false negative problem in existing image-text matching benchmarks, [13] proposed an ITM framework integrating Language Guidance. In terms of improving retrieval efficiency and accuracy, [14] proposed a self-supervised fine-grained alignment module called SelfAlign. [15] proposed a feature-first approach to advance image-text retrieval by improving visual features. This method generates more comprehensive and robust visual features through a two-step interaction model and a multi-attention pooling module. Recently, with the development of large-scale pretraining models like CLIP [16] and ALIGN [17], significant performance improvements are achieved in image-text matching tasks by pretraining on massive vision-language data and then fine-tuning for specific downstream tasks. Different from the image retrieval methods, we focus on retrieving the image based on the conversation not just the text.
@@ II-C Multi-modal Conversation
Several multi-modal studies have sought to improve the efficacy of conversational agents by enriching textual expressions of generated dialog responses through associative vision scenes [18, 19, 20, 21]. In contrast, sticker retrieval in the real-world requires understanding the semantic expression and intention of the context and the sticker image. Several works specifically paid attention to sticker features [22, 4]. [22] proposed a real-time system for sticker recommendation, which decomposes the sticker recommendation task into two steps, including message prediction that the user is likely to send and an appropriate sticker substitution. Nowadays, using stickers for replies has become commonplace in social media interactions. There has been a growing number of works on sticker retrieval, which assists users in selecting the appropriate sticker for a response. [8] proposed a sticker response selector model that employs a convolution-based sticker image encoder and a self-attention-based multi-turn dialog encoder to obtain the representation of stickers and utterances. Then, a deep interaction network is designed for obtaining deep matching, and a fusion network is employed to output the final matching score. [7] presented a Meme incorporated Open-domain Dialogue (MOD) task and utilized pooling all sub-tasks like text generation and internet meme prediction into a unified sequence generation procedure to solve it. To tackle three challenges of inherent multimodality, significant inter-series variations, and complex multi-modal sentiment fusion. [4] integrated both the sticker text and series tag to holistically model sticker sentiment. They employed a flexible masked attention mechanism to selectively extract the most pertinent information crucial for the current sentiment assessment. These methods, which directly match conversation context and stickers, overlook the expressive role of stickers in conversation, where emotions and intentions are conveyed in a visually engaging manner.
@@ III StickerInt Dataset
This section introduces our new dataset StickerInt for sticker retrieval in detail. Specifically, Section III-A presents the construction of our StickerInt dataset, Section III-B introduces the annotation process of StickerInt, and Section III-C analyzes the created StickerInt by data statistics.
@@ III-C Dataset Analysis
The detailed statistics of our dataset are shown in Table I. In total, there are 1,578 conversations that contain 12,644 utterances, 59,623 tokens, and 1,025 stickers. Of these stickers, 702 are associated with textual content. The proposed dataset is collected from conversations with Chinese users on WeChat, and the dialogues are predominantly in Chinese. Notably, 841 utterances include English words such as “OK”, “QQ”, and “NLP”, which are naturally embedded in Chinese expressions, reflecting the way people commonly communicate in daily life. Each conversation includes 8.01 utterances on average. The average number of users who participate in a conversation is 2.80. Among all conversations in our dataset, 1,017 belong to the Supplementary Response scenario, and 561 fall under the Direct Response scenario (sticker as a direct reply to previous conversation turns). The average number of tokens in an utterance is 5.21. Furthermore, we also visualize the distribution of intention tags in Figure 2. We can see that the proportion of stickers used in different labels varies, which also demonstrates the diversity of stickers’ intention expressions in our dataset.
@@ IV Methodology
In this section, we introduce our proposed Int-RA framework for sticker retrieval in detail. Assume that there is a conversation context $T=\{t_{1},...,t_{N_{t}},s_{j}\}$ and a sticker set $V=\{v_{1},...v_{N_{v}}\}$, where $t_{i}=(s_{i},u_{i})$, $s_{i}$ and $u_{i}$ represent the $i$-th user and utterance. $N_{t}$ and $N_{v}$ represent the number of utterances and stickers, respectively. The sticker retrieval task aims to select a suitable sticker from the sticker set $V$ based on the conversation context $T$ for the user $s_{j}$. In the $i$-th utterance $u_{i}=\{w_{1}^{i},...,w^{i}_{N_{w}^{i}}\}$, $w_{j}^{i}$ represents the $j$-th word in $u_{i}$, and $N_{x}^{i}$ represents the total number of words in $u_{i}$ utterance. Additionally, our dataset presents an annotated intention tag $y_{int}$ as supplemental information in each conversation $T$. In the dialog context $T$, $s_{i}$ represents a sticker image with a binary label $y_{i}$, indicating whether $s_{i}$ is an appropriate response for $T$. Each sticker with the intention label $m_{i}$ indicates the intention of the speaker to use the sticker to respond. Our goal is to learn a ranking model that can produce the correct ranking for each dialog.
@@ V-C Main Results
For text-to-image retrieval methods, they prioritize capturing semantic relationships between text and image content. However, since they are not explicitly designed for sticker retrieval scenarios, exhibiting inferior performance compared to our framework. In addition, an interesting observation is the overall superior performance of text-to-image retrieval methods compared to sticker retrieval baseline methods. This disparity can be attributed in part to the model design of SRS and MOD, where both models are devised to retrieve suitable stickers from a limited set of similar sticker candidates. Consequently, more attention is devoted to distinguishing local information among similar sticker expressions. In contrast, our dataset encompasses a more diverse range of stickers in real-world conversations, requiring the model to analyze and recognize more complex visual features. This demonstrates the advantage of the new dataset and the proposed framework in this work.
For large language models, except for Ultron and ChatIR, we design a prompt that integrates the current session to generate the sticker description for each session, leveraging the in-context learning ability of LLMs [49, 50]. We then retrieve the appropriate sticker based on the generated sticker description and the combination of intent and sticker attributes using OpenAI’s LLM-based embedding model (text-embedding-ada-003) [51]. We can observe that LLM models perform better than sticker retrieval and text-to-image retrieval methods. This superior performance is attributed to the extensive parameterization and complex network architecture of LLMs, which greatly enhance their ability to understand and generate complex language and image descriptions. For P@1, Ultron performs the best, while for P@3 and P@5, ChatGLM3 achieves the highest performance, reaching 25.32% and 31.17%. Compared to our IGSR method, all baseline models perform significantly worse. This further emphasizes the effectiveness of our approach in intent derivation, highlighting the critical role of intent as a bridging component in the process.
@@ V-E Ablation Study
We also conduct an ablation study on the use of knowledge and attributes. The evaluation results are shown in Table IV. The performances of all ablation models are worse than those of the complete model under all metrics, which demonstrates the necessity of each component in our approach. Note that the removal of attributes ("w/o attribute") results in considerable performance degradation, indicating that utilizing attributes can lead to better learning of sticker representation in different sticker properties. Notably, by observing the performance of w/o attribute on P@5, we find that the impact of attributes is not as significant when a larger number of stickers are recalled. In addition, the removal of intention tag ( "w/o intention" ) sharply degrades performance, which verifies the importance of knowledge in understanding conversation context. This demonstrates the crucial role of the intention tag in improving the accuracy and relevance of sticker retrieval tasks, highlighting the importance of considering users’ intentions in retrieving appropriate stickers.
@@ V-G Effect of Number of Utterances
To examine and analyze the impact of the number of utterances on the performance of our proposed Int-RA framework, we conduct experiments by varying the number from 2 to 10 and demonstrate the results in Figure 6. We observe a similar trend across all evaluation metrics: mAP, P@1, P@3, and P@5. The results initially increase until the number of utterances reaches 6, after which they decrease as the number of utterances continues to increase. One potential reason may explain this phenomenon. As the number of utterances increases from one to six, the model benefits from richer contextual information, which enhances its understanding of the conversation and improves retrieval performance. However, when the context becomes too long, utterances that appear much earlier than the sticker response may become less relevant, introducing noise and hindering accurate sticker selection. In our dataset, using six utterances offers the best balance between context richness and noise, leading to optimal performance.
@@ V-K Visualization
To analyze how our Int-RA learns the important information about stickers, we visualize the relation score $\mathcal{M}$ (Equation 12) of three stickers in Figure 8. For example (a), where the character appears neutral. This indicates that the representation of this sticker heavily relies on this facial expression. Our Int-RA can effectively catch the important information in the sticker by the relation score placement on the character’s face. Moreover, the relation score also attends to the character’s gestures. For instance, in Case (b), where the character is depicted as holding chopsticks with one hand and supporting the face with the other, we observe attention focused on his hand, suggesting that our Int-RA learns key points of body language. Furthermore, considering that the relation score comprehensively considers four properties of stickers, as illustrated in Case (c), we observe that our Int-RA pays attention to both facial expressions and gestures simultaneously, thereby learning accurate visual information about the sticker.
@@ Acknowledgments
This work was partially supported by the National Natural Science Foundation of China (62176076), the Natural Science Foundation of Guangdong (2023A1515012922), the Shenzhen Foundational Research Funding (JCYJ20220818102415032), the Guangdong Provincial Key Laboratory of Novel Security Intelligence Technologies (2022B1212010005), Hong Kong RGC GRF No. 14206324, CUHK direct grant No. 4055209, and CUHK Knowledge Transfer Project Fund No. KPF23GWP20.